n a certain material that has index of refraction of 2.25, and electromagnetic wave has speed of 7.5 x 107 m/s, and a wavelength of 6.4 x 10-9 m. Calculate the frequency of the EM wave when it is propagating in vacuum.
Answer #1

Frequency of wave is not change with medium,it is always same in each medium and vacuum.

We know f= speed / wavelength

Speed =7.5× 10^7 m/s

Wavelength = 6.4× 10^(-9) = 6.4nm

So frequency =1.172× 10^(16) Hz